Afirs production engineers have estimated its production function as Q(KL) = 2K2/1/2 = 2-K L=2KL where and L represent t
Answer #1

Cost minimization condition,

MPk/r=MPL/w

√L/r√K=√K/w√L

w*L=rK

L=100k/16=25k/4

Q=10

10=2√25k^2/4=2*5K/2=5K

K=10/5=2

L=25*2/4=12.5

C=12.5*16+2*100=400

Q=30

30=5K

K=6

L=25*6/4=37.5

C=37.5*16+6*100=1200

So given Q=30 ,C=400

Q=30 ,C=1200

So we can derive cost function,C=40 Q

Option b is correct